OpenStudy (anonymous):

You are buying gifts for 10 people. You decide to buy each person either a CD or a DVD. A CD costs $12 and a DVD costs $20. a. Let c=the number of CDs you decide to buy. What is an expression in terms of c for the number of DVDs you buy? b. What is an expression in terms of c for the cost of the CDs? For the cost of the DVDs? c. Write and simplify an expression in terms of c for the total cost of all the gifts you buy. What properties of real numbers did you use to simplify the expression.

OpenStudy (valpey):

You just have to write it out like a sentence: "The total cost is the number of DVDs I buy times $20 plus the number of CDs I buy times $12. The total number of gifts I buy is 10. This is equal to the number of DVDs I buy plus the number of CDs I buy" Now transform the sentence into an equation using the following replacements: c = "number of CDs I buy" d = "number of DVDs I buy" t = "The total cost" "The total cost is the number of DVDs I buy times $20 plus the number of CDs I buy times $12. The total number of gifts I buy is 10. This is equal to the number of DVDs I buy plus the number of CDs I buy" becomes: "t is the d times $20 plus c times $12. The total number of gifts I buy is 10. This is equal to the d plus c" Now let's replace words like equals, plus, times... t = d*$20 + c*$12 10= d + c

OpenStudy (valpey):

Now we can solve part a.) What is an expression in terms of c for the number of DVDs you buy?

OpenStudy (valpey):

We just need the second equation to do this, so starting with: 10 = d + c ...

OpenStudy (anonymous):

so would it be 10d=c?

OpenStudy (anonymous):

What is it ??? I'm on it right now